OpenStudy (imamod):

FAN + MEDAL If the cos of angle x is 8 over 17 and the triangle was dilated to be two times as big as the original, what would be the value of the cos of x for the dilated triangle? Hint: Use the slash symbol ( / ) to represent the fraction bar, and enter the fraction with no spaces.

OpenStudy (solomonzelman):

the cosine of an angle doesn't change, because the angle doesn't change. In other words, 2/4 is the same thing as 1/2.

OpenStudy (imamod):

can you show me ?

OpenStudy (imamod):

@SolomonZelman

OpenStudy (imamod):

@satellite73

OpenStudy (mokeira):

@SolomonZelman Well said

OpenStudy (imamod):

@Mokeira ?

OpenStudy (imamod):

can you help

OpenStudy (mokeira):

Let me try and illustrate

OpenStudy (mokeira):

uuuuh...my network has refreshed. I have to start all over

OpenStudy (mokeira):

You have the cos so we can sketch the triangle  When it is doubled it becomes  Now, get cos x of this new triangle \[\cos x=\frac{ 16 }{ 34 }\] This fraction can be simplified \[\frac{ 8 }{ 17 }\]

OpenStudy (imamod):

thanks

OpenStudy (mokeira):

my network jammed...sorry for taking tooooooo long